OpenStudy (anonymous):

f y varies inversely as x, and y = 3 as x = -2, find y for the x-value of -1. please help :(

OpenStudy (psymon):

Varies inversely means you use the formula: \[y=\frac{ k }{ x } \]where k is some unkwnon constant we have to solve for. So based on what is given, it says y = 3 and x = -2. Using these two numbers, we can plug them into the formula and solve for k. \[3=\frac{ k }{ -2 }\]Multiplying both sides by -2, we get that k = -6. Now this value of k can be used to solve the last part of the problem. So we need to find y when x = -1. Well, knowing that x = -1 and k = -6, we can plugthose into the formula as well to get: \[y=\frac{ -6 }{ -1 }\]therefore y = 6. Make sense?

OpenStudy (anonymous):

wow man thx soo much!

OpenStudy (psymon):

Yeah, np ^_^

OpenStudy (anonymous):

one more If y varies directly as x, and y = 15 as x = -3, find y for the x-value of 4. i have 20 more to do and this one is a bit different

OpenStudy (anonymous):

im doing the other ones if youre thinking there is no way on the earth im going to do 20 lol

OpenStudy (psymon):

Well, if its directly (can also say jointly, means the same thing), the process is the same but the formula is different. Instead, we have this formula: \[y=kx \] So basically k times x instead of k divided by x. So same thing, we use the first numbers to get k. So if y = 15 and x = -3, we just plug those into the formula to get: 15 = -3k. Divide by -3 and we get k = -5. So using k = -5, we use that to help us solve for y with the other numbers. So if x = 4, and we know k = -5, we can say y = (-4)(5), meaning y = -20. And if you gert the hang of these then 20 wont be super bad.

OpenStudy (anonymous):

no it wont be bad :P thx for your help i got the hang of it

OpenStudy (psymon):

Awesome : ). One last thing, though. It wont always say y varies. It might say something like the height of a person varies inversely as the square of the person's weight. Or something weird like that. Just know that the phrase _______ varies. Whatever that blank is is what your y is in the formula. The other vaue or information given is x. So for what problem I just invented, it would be \[y = \frac{ k }{ x } \] \[height = \frac{ k }{ (weight)^{2} } \]Just wanted to show ya an example of that.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."